### v2

# LLM-Driven Multi-step Translation from C to Rust using Static Analysis

Abstract
Translating software written in legacy languages to modern languages, such as C to Rust, has significant benefits in improving memory safety while maintaining high performance. However, manual translation is cumbersome, error-prone, and produces unidiomatic code. Large language models (LLMs) have demonstrated promise in producing idiomatic translations, but offer no correctness guarantees as they lack the ability to capture all the semantics differences between the source and target languages. To resolve this issue, we propose SACTOR, an LLM-driven C-to-Rust zero-shot translation tool using a two-step translation methodology: an “unidiomatic” step to translate C into Rust while preserving semantics, and an “idiomatic” step to refine the code to follow Rust’s semantic standards. SACTOR utilizes information provided by static analysis of the source C program to address challenges such as pointer semantics and dependency resolution. To validate the correctness of the translated result from each step, we use end-to-end testing via the foreign function interface to embed our translated code segment into the original code. We evaluate the translation of $200$ programs from two datasets and two case studies, comparing the performance of GPT-4o, Claude 3.5 Sonnet, Gemini 2.0 Flash, Llama 3.3 70B and DeepSeek-R1 in SACTOR. Our results demonstrate that SACTOR achieves high correctness and improved idiomaticity, with the best-performing model (DeepSeek-R1) reaching 93% and (GPT-4o, Claude 3.5, DeepSeek-R1) reaching 84% correctness (on each dataset, respectively), while producing more natural and Rust-compliant translations compared to existing methods.
Keywords Software Engineering  $\cdot$ Static Analysis  $\cdot$ C  $\cdot$ Rust  $\cdot$ Large Language Models  $\cdot$ Machine Learning
1 Introduction
The C language is the most widely used language in system-level programming. C’s ability to directly manipulate memory and hardware has made it indispensable in developing operating systems, embedded systems, and performance-critical applications (Love, 2013). However, its manual memory management has led to memory-related bugs, such as buffer overflows, dangling pointers, and memory leaks, which have been the root cause of many security vulnerabilities and system crashes (Fan et al., 2020).
To address these shortcomings, Rust has emerged as a promising alternative to C. Rust is a modern system programming language that offers memory safety without relying on garbage collection by imposing a strict object ownership model (Matsakis and Klock, 2014). Due to Rust’s memory safety guarantees and performance, it is adopted by projects like the Linux kernel in developing new hardware drivers [^1] https://github.com/Rust-for-Linux/linux and Mozilla Firefox. Translating legacy C code to Rust provides a path to improve safety and performance, but manual translation is time-intensive, error-prone, and requires developer expertise in both languages.
Researchers have started investigating automatic translation methods for C-to-Rust, resulting in tools such as C2Rust (Immunant, 2020) which analyzes the abstract syntax tree (AST) of the code to produce syntactically equivalent Rust code. Such “traditional” approaches for C-to-Rust translation often rely on rule-based techniques or static analysis (Zhang et al., 2023; Immunant, 2020; Emre et al., 2021; Hong and Ryu, 2024; Ling et al., 2022). While these methods are effective at producing syntactically correct translations, they often fail to provide an “idiomatic” translation: one that follows the Rust programming model and conventions with the least use of unsafe blocks. Due to fundamental differences in semantics and memory management between C and Rust, idiomatic translations are crucial to compiler-enforced memory safety, as well as improving code readability and maintainability.
Large language models (LLMs) have demonstrated the ability to understand syntax and code semantics, making them a promising tool for code translation (Pan et al., 2023). However, LLMs are prone to hallucinations and often generate incorrect code (Perry et al., 2023). Moreover, the large semantic differences between C and Rust make it more challenging for LLMs to capture correct semantics of the code, leading to inaccurate translations. Previous work has explored various verification techniques, such as fuzz tests and symbolic execution (Yang et al., 2024; Eniser et al., 2024) to validate LLM-generated code. Although these methods can guarantee equivalence, their strictness often limits scalability or applicability to code with complicated code features. For example, Vert (Yang et al., 2024) always struggles to verify the correctness of code with data structures through symbolic execution.
In this paper, we introduce Structure-Aware C-to-Rust Translator (or SACTOR), an LLM-driven zero-shot translation tool. An overview of SACTOR is presented in Figure 2. We divide the translation process into two steps: 1. C to unidiomatic Rust where we focus on syntactic translation only and allow C-like semantics and unsafe code blocks to keep all the unsafe memory manipulations, and 2. unidiomatic Rust to idiomatic Rust where we refine the translation to follow Rust semantic standards and remove all the unsafe blocks in the code to ensure safety. In both translation steps, we leverage static analysis tools to provide critical information on pointers and code dependencies and guide the translation process. To address the challenge of verifying LLM outputs efficiently, we utilize end-to-end testing by embedding the translated Rust code with the original C code through foreign function interface (FFI). This allows us to directly run the end-to-end tests on the whole program to validate a particular translation. This combination of static analysis, multi-step refinement, and efficient (and robust) verification significantly improves translation quality, achieves high-level correctness, and provides a broader applicability to different types of C programs.
We make the following contributions:
  • We propose SACTOR, a tool that integrates feedback from static analysis tools into a two-step translation process. By doing so, we reliably obtain idiomatic unsafe-free translated Rust from C (§ 4).
  • We propose and demonstrate a new verification approach by embedding the translated code into the original code and utilizing the end-to-end testing. This enables more practical, scalable, and efficient verification method for LLM-generated code (§ 4.3).
  • We evaluate our approach on a diverse set of C programs across two datasets and five different LLMs. Our best-performing model, DeepSeek-R1 achieves 93% success rate on the first dataset, while GPT-4o, Claude 3.5, and DeepSeek-R1 reach 84% on the second dataset (§ 6.1.1). Notably, SACTOR generates more idiomatic, unsafe-free Rust code compared to four existing methods (§ 6.3). These results demonstrate the broad applicability and effectiveness of SACTOR. [^2] SACTOR code is available at https://github.com/qsdrqs/sactor and datasets are available at https://github.com/qsdrqs/sactor-datasets.
  • We compare the cost of each LLM in terms of token consumption and average queries required for a single translation and observe that GPT-4o is the most efficient in terms of token consumption while DeepSeek-R1 consumes 5-7$\times$ more tokens than GPT-4o. On the other hand, We observe that the average number of queries varies across models but is less pronounced than token consumption. The most query-intensive model, Llama 3.3, requires only about 1.4$\times$ more queries than the least costly model, Gemini 2.0 (§ 6.2).
  • Finally, we inspect failure cases and find that different LLMs exhibit varying strengths and weaknesses in translating C code to Rust (§ 6.1.2). DeepSeek-R1 excels in handling format string errors, array manipulation, and has a lower failure rate in different categories of errors on the TransCoder-IR dataset, which features complex C code. This suggests reasoning models like DeepSeek-R1 perform better in translating code with intricate logic and data manipulations.
2 Background
2.1 Primer on C and Rust
C is one of the most widely used programming languages (TIOBE, 2025). It is a low-level language that provides direct access to memory and hardware through pointers and abstracts the machine-level instructions. While this makes it efficient, it suffers from memory vulnerabilities such as buffer overflow (MITRE, 2006a, b), dangling pointers (MITRE, 2006c), and memory leaks (MITRE, 2006d).
Rust, in contrast, is a modern programming language that provides memory safety without additional performance penalty, and has the same ability to access low-level hardware as C. Rust achieves this by enforcing strict rules at compile time such as (1) ownership, which allows only one owner for each piece of memory; (2) borrowing, which allows multiple immutable references or one mutable reference to the memory; and (3) lifetimes, which track the scope of the variables to ensure their memory is freed when they no longer used (Matsakis and Klock, 2014).
2.2 Challenges in Code Translation
C Code:
Rust Code:
Figure 1: Example of a simple C program and its equivalent Rust program, both hand-written for illustration.
Why do we need to translate C to Rust? Despite its features, Rust is a relatively new language; there exist many widely-used system-level programs written in C, but its lack of memory safety leads to vulnerabilities (§ 2.1). It is desirable for these programs to directly benefit from Rust’s memory safety features and enhanced security and reliability.
However, translating C to Rust is challenging due to fundamental differences between the two languages. Figure 1 shows an example of a simple C program and its Rust equivalent. The create_sequence function takes an integer n as input and returns an array with a sequence of integers. In C, the function needs to allocate memory for the array using malloc and will return the pointer to the allocated memory as an array. If the size is invalid, or the allocation fails, the function will return NULL. The caller of the function is responsible for freeing the memory using free when it is done with the array to prevent memory leaks.
In contrast, the Rust implementation uses Vec<i32> with automatic memory management. If the size is invalid, the function returns None using the Option type (an idiomatic choice). The caller does not need to free the memory, as the Vec will be automatically de-allocated when it goes out of scope. Although Rust allows the use of unsafe blocks to enable operations like pointer manipulation similar to C, it is discouraged due to the lack of safety guarantees from compiler and is considered unidiomatic.
There are other differences between C and Rust, such as representation of strings, use of pointers, handling of arrays, lifetime of references, and the error handling mechanism. A non-exhaustive list of differences between C and Rust is shown in Appendix A.
All these differences make the translation process challenging. An ideal translator must understand (1) the semantics of both languages, (2) the context, logic and algorithms of the C code to correctly handle the memory management differences, and (3) other language-specific features in order to correctly translate. Additionally, we would also require a practical way to verify the correctness of the translated code.
3 Related Work
We believe that large language models (LLMs) can act as ideal translators for this particular task. State-of-the-art LLMs are trained on large volumes of code (OpenAI, 2022), and are expected to be aware of language-specific constructs. Additionally, programs are often provided with comments, enabling LLMs to understand fine-grained details needed for successful translation.
3.1 LLMs for C-to-Rust Translation
Vert (Yang et al., 2024) uses LLMs to generate candidate translated (Rust) code and then adapts fuzz testing and symbolic execution to prove the equivalence. This provides a strong equivalence from the syntactic and semantic perspective. However, its symbolic execution is limited in scalability and struggles with complex C features. Eniser et al. (2024) proposed Flourine, which utilizes error messages as feedback to the LLMs to guide translation. They use fuzz testing and data type serialization to tackle the type mismatch problem. However, data type serialization issues account for half of their translation errors. Shiraishi and Shinagawa (2024) segment C code into smaller sub-tasks to overcome LLMs’ context-length limitations. They directly guide LLMs to translate specific language features, such as macros, using predefined Rust conventions. However, they only measure compilation success, without ensuring functional equivalence. Shetty et al. (2024) use dynamic analysis to analyze the runtime behavior of the C code, and uses this information to guide LLM translation. However, dynamic analysis is limited by input coverage, making it challenging to generalize across all execution paths. Nitin et al. (2025) refines C2Rust’s translation instead of directly converting C to Rust using an LLM. It employs an LLM to reduce unidiomatic Rust code, particularly unsafe blocks and libc functions. However, the output still contains a considerable amount of unidiomatic Rust and does not substantially reduce the reliance on unsafe blocks and libc functions. Furthermore, the translation quality is limited by C2Rust itself. As noted in § 4.2.1, C2Rust diminishes code readability and strips comments and pre-processor directives from the original C code, making it harder for the LLM to interpret the code and produce an accurate translation.
3.2 Non-LLM Approaches for C-to-Rust Translation
C2Rust (Immunant, 2020) translates C-to-Rust by extracting the abstract syntax tree (AST) of the C code and transforming the C AST to the corresponding Rust AST. It then uses a rule-based translator to convert the Rust AST to Rust code. However, the output is a direct syntactic translation, and relies entirely on unsafe blocks. Additionally, using a rule-based coverter to generate Rust code from AST will significantly reduces readability because it directly translates C constructs into Rust without abstraction and excessive explicit type conversions. Crown (Zhang et al., 2023) is a tool that can analyze the pointer information in C code based on its static ownership tracking. It uses this information to generate Rust code with the corresponding semantics and reduce the pointer usage in the translated code. Hong and Ryu (2024) proposed an approach to handle the return value of functions in C-to-Rust translation. Ling et al. (2022) translate Rust based on a set of predefined rules and heuristics. Although these approaches reduce the use of unsafe blocks compared to C2Rust, the translated code still heavily relies on unsafe blocks and remains largely unidiomatic.
4 SACTOR Methodology
We propose SACTOR, an LLM-driven C-to-Rust translation tool using a two-step translation methodology. The overview of SACTOR’s methodology is shown in Figure 2. Recall that the semantics of Rust is different from that of C (§ 2). To assist the LLM overcome this gap, and capture more nuanced semantic information about the C code, we utilize a suite of static analysis tools to provide additional information in the form of hints to the LLM. To this end, we outline the four main stages below.
Figure 2: Overview of the SACTOR methodology.
4.1 Task Division
We begin by dividing the program into smaller parts that can be processed by the LLM independently. This helps the LLM focus on a narrower scope for each translation task, and ensures that our program fits within the LLM’s context window. To facilitate task division and extract relevant language information—-such as definitions, declarations, and dependencies—from C code, we developed a static analysis tool called C Parser based on libclang. libclang is a library that provides a C compiler interface, allowing access to semantic information of the code. Our C Parser analyzes the input program and splits the program into fragments consisting of a single type, global variable, or function definition. This step also extracts semantic dependencies between each part (e.g., a function definition depending on a prior type definition). We then process each program fragment in dependency order: all dependencies of a code fragment are processed before the fragment. In Appendix B, we explain how we break up the program into pieces in more detail.
4.2 Code Translation
To ensure that each program fragment is translated only after its dependencies have been processed, we begin by translating data types, as they form the foundational elements for functions. This is followed by global variables and functions, for both unidiomatic translation and idiomatic translation. We divide the translation process into two steps.
4.2.1 Unidiomatic Rust Translation
We aim to produce semantically equivalent Rust code from the original C code, allowing the use of unsafe blocks and C standard library functions.
For data type translation, we leverage the C2Rust tool (Immunant, 2020) to convert C code into Rust code. While C2Rust provides reliable data type translation, it struggles with function translation due to its compiler-based approach, which omits source-level details like comments, macros, and other elements. These omissions significantly reduce the readability and usability of the generated Rust code. Thus, we use C2Rust only for data type translation and use an LLM to translate global variables and functions. For functions, we rely on our C parser to automatically extract dependencies (e.g., function signatures, data types, and global variables) and reference the corresponding Rust code. This approach guides the LLM to accurately translate functions by leveraging the previously translated components and directly reusing or invoking them as needed.
4.2.2 Idiomatic Rust Translation
The goal of this step is to refine unidiomatic Rust code into idiomatic Rust by removing unsafe blocks and following Rust idioms.
Handling pointers from C code is a key challenge, as they are considered unsafe in Rust. Unsafe pointers should be replaced with Rust types such as references, arrays, or owned types. To address this, we use Crown (Zhang et al., 2023) to facilitate the translation by analyzing pointer mutability, fatness (e.g., arrays), and ownership. This information provided by Crown helps the LLM assign appropriate Rust types to pointers. Owned pointers are translated to Box, while borrowed pointers use references or smart pointers. Crown assists in translating data types like structs and unions, which are processed first as they are often dependencies for functions. For function translations, Crown analyzes parameters and return pointers, while local variable pointers are inferred by the LLM. Dependencies are extracted using our C parser to guide accurate function translation.
4.3 Verification
To verify the equivalence between source and target languages, prior work has relied on symbolic execution and fuzz testing. However, these methods are impractical for real-world C-to-Rust translation (details in Appendix C). We propose a new approach to validate the correctness of translated Rust code by focusing on soft equivalence—ensuring functional equivalence of the entire program based on success on end-to-end tests. This approach avoids the complexity of generating specific inputs or constraints for individual functions and is well-suited for real-world programs where such end-to-end tests are often available and reusable.
4.3.1 Verifying unidiomatic Rust code
Verifying the unidiomatic Rust code is straightforward; it is semantically equivalent to the C code, with compatible function definitions and data types. First, the unidiomatic Rust code is compiled using the Rust compiler to check for successful compilation. Once it is compiled, the original C code will be recompiled by linking it to the unidiomatic Rust translation as a shared library. This setup ensures that when the C code calls the target function, it invokes the Rust translation instead. To verify correctness, end-to-end tests are run on the entire program, comparing the outputs of the original C code and the unidiomatic Rust translation. If all tests pass, the target function is considered verified.
4.3.2 Verifying idiomatic Rust code
The verification of idiomatic Rust code is more complex, as the idiomatic Rust code does not align with the original C code in data types or function definitions. Direct linking between the original C code and the idiomatic Rust code is therefore infeasible. To address this, we use the LLM to create a test harness for each target function. The test harness mirrors the function definition of the unidiomatic Rust version, allowing it to accept C data types as input and return C data types as output. Within the harness, inputs are converted from C data types to their corresponding Rust types before calling the target function. The output is then converted back from Rust types to C types to maintain type consistency. Figure 3 provides an example, where a string and an integer are passed as input, and an owned string is returned. Input strings are converted from C’s char* to Rust’s &str, and output strings are converted from Rust’s String back to C’s char*.
Figure 3: Example of the test harness function.
Type conversion in the harness involves two cases:
  • Basic types: Simple conversions like int to i32, float to f32, and char* to String or &str, which are straightforward for the LLM to handle.
  • Custom data structures: These conversions are more complex due to layout differences between C and Rust structures. For such cases, the LLM firstly generates two conversion functions for each structure—one for converting from C-to-Rust and another for converting from Rust-to-C. These functions are tailored to specific data structures and are invoked by the test harness.
Once the harness is constructed, it can be linked to the original C code, as in § 4.3.1. End-to-end tests are then executed for the entire program, enabling verification of the idiomatic Rust code even when it diverges from the original C code in data types and function definitions.
4.3.3 Feedback Mechanism
For code that fails the verification process, feedback is provided to the translation process to help the LLM correct the issues. If the Rust code fails to compile, the compiler errors are directly passed back to the translation process to guide the LLM in fixing the translation. For end-to-end test failures, a Rust procedural macro is employed to insert debugging code into the target function during compile time. This debugging code logs the inputs and outputs of the target function when a test fails. The end-to-end tests are then re-executed, and the collected information is fed back to the translation process to assist the LLM in resolving the errors.
4.4 Code Combination
By translating and verifying all functions and data types, we can integrate them into a unified Rust code-base. We first gather the translated Rust code from each sub-task, then eliminate duplicate use statements and other redundancies needed for individual sub-task compilation. Next, we organize the code into a well-structured, idiomatic implementation of the original C program. Once combined, the entire program is subjected to end-to-end tests to verify the correctness of the final Rust code. If all tests pass, the translation process is deemed successful.
5 Experimental Setup
5.1 Datasets Used
Table 1: Summary of datasets and real-world projects used for evaluation.
Dataset Size Preprocessing E2E Tests Corresponding Rust
TransCoder-IR (Szafraniec et al., 2023) 100 Removed buggy programs (compilation and memory errors) and programs that have Rust translation Present Absent
Project CodeNet (Puri et al., 2021) 100 Filtered for programs with external input (argc/argv) Absent Absent
Real-World Projects 2 Extend macros, combine the whole project to a single file Present (Limited) Absent
For the selection of datasets for evaluation, we consider the following criteria: (1) Absence of Corresponding Rust Code: The dataset must not include equivalent Rust code to prevent LLMs from simply memorizing and reproducing the dataset instead of performing genuine translation; (2) Sufficient Number of C Programs: The dataset should contain a substantial number of C programs to ensure a robust evaluation of the approach’s performance across a diverse set of examples; (3) Presence of Non-Trivial C Features: The dataset should include C programs with advanced features such as multiple functions, structs, and other non-trivial constructs as it enables the evaluation to assess the approach’s ability to handle complex features of C; and (4) Availability of End-to-End Tests: The dataset should either include end-to-end tests or make it easy to generate them as this is essential for accurately evaluating the correctness of the translated code through comprehensive testing.
Based on the above aspects, we select two datasets and two real-world projects. Their salient details are presented in Table 1, and complete details are presented in Appendix D.
5.2 Evaluation Metrics
Success Rate: The success rate is defined as the ratio of the number of programs that can (a) successfully be translated to Rust and (b) successfully pass the end-to-end tests for both unidiomatic and idiomatic translation phases to the total number if programs. To enable the LLMs to utilize feedback from previous failed attempts, we allow the LLM to make up to 6 attempts for each translation process (each using feedback from all previous attempts). After all 6 attempts, we treat this test case as failure.
Idiomaticity: To evaluate the idiomaticity of the translated code, we use two metrics.
1. Lint Alert Count is measured by running Rust-Clippy (The Rust Team, 2016), a tool that provides lints on unidiomatic Rust code (including improper use of unsafe code and other common style issues). By collecting the warnings and errors generated by Rust-Clippy for the translated code, we can assess its idiomaticity: fewer alerts indicate more idiomatic translation. Previous works (Yang et al., 2024; Eniser et al., 2024) have also used Rust-Clippy for similar evaluations.
2. Unsafe Code Fraction, inspired by Shiraishi and Shinagawa (2024), is defined as the ratio of tokens inside unsafe code blocks or functions to total tokens for a single program. High usage of unsafe is considered unidiomatic in Rust, as it bypasses compiler safety checks, introduces potential memory safety issues and reduces code readability.
3. Unsafe Free Fraction indicates the percentage of translated programs in a dataset that do not contain any unsafe code. Since unsafe code represents potential points where the compiler cannot guarantee safety, this metric helps determine the fraction of results that can be achieved without relying on unsafe code.
Cost: To evaluate the cost of our approach, we measure (1) Total LLM Queries as the number of total LLM queries made during translation and verification for a single test case in each dataset , and (2) Total Token Count as the total number of tokens processed by the LLM for a single test case in each dataset. To ensure a fair comparison across models, we use the same tokenizer (tiktoken) and encoding (o200k_base).
5.3 LLMs Used
We evaluate our approach using GPT-4o (OpenAI), Claude 3.5 Sonnet (Anthropic), Gemini 2.0 Flash (Google), DeepSeek-R1 (DeepSeek) and Llama 3.3 70B Instruct (Meta)—a fine-tuned variant of Llama 3.3 70B optimized for text generation. Except for DeepSeek-R1, all models are non-reasoning models, i.e. directly generate output without chain-of-thought reasoning. LLM configurations are detailed in Appendix E.
6 Evaluation
Through our evaluation, we aim to answer the following research questions: (1) How successful is SACTOR in generating idiomatic Rust code using different LLM models?; (2) What are the computational costs associated with SACTOR?; (3) How idiomatic is the Rust code produced by SACTOR compared to existing approaches?; and (4) How well does SACTOR generalize to complex code-bases?
Our results show that: (1) DeepSeek-R1 achieves the highest success rates (93%) with SACTOR on TransCoder-IR and 3 models (GPT-4o, Claude 3.5, DeepSeek-R1) reaches the highest success rates (84%) on Project CodeNet, while failure reasons vary between datasets and models (§ 6.1); (2) GPT-4o and Gemini 2.0 offer the best balance between performance and cost, whereas Llama 3.3 had the highest token consumption among non-reasoning models. DeepSeek-R1 required approximately 3 to 7 times more tokens compared to other non-reasoning models (§ 6.2) ; (3) SACTOR’s idiomatic translation results outperforms all previous baselines, producing Rust code with fewer Clippy warnings and 100% unsafe-free translations (§ 6.3); and (4) SACTOR successfully generates the verified unidiomatic translation results on complex code-bases like avl_tree and urlparser, but has difficulties verifying idiomatic translations when handling function pointers and complex data conversions (§ 6.4).
6.1 Generating Idiomatic Rust Code
(a) TransCoder-IR SR
(b) CodeNet SR
Figure 4: Success rates (SR) across different LLM models for the TransCoder-IR and CodeNet datasets. SR 1-6 represent the number of attempts made to achieve a successful translation in idiomatic translation.
6.1.1 Success Rate Evaluation
We evaluate the success rate (as defined in § 5.2) for two datasets on different models. For idiomatic translation, we also plot how many attempts are needed.
(1) TransCoder-IR (Figure 4(a)): DeepSeek-R1 achieves the highest success rate (SR) in both unidiomatic (94%) and idiomatic (93%) steps, only 1% drop in the idiomatic translation step, demonstrating strong consistency in code translation. GPT-4o follows with 83% in the unidiomatic step and 78% in the idiomatic step. Gemini 2.0 comes next with 78% and 75%, respectively. Claude 3.5 struggles in the unidiomatic step (55%) but does not show substantial degradation when converting unidiomatic Rust to idiomatic Rust (54%, only a 1% drop). Llama 3.3 performs better in the unidiomatic step (68%) but drops significantly in the idiomatic step (50%), and requiring more attempts for correctness.
(2) Project CodeNet (Figure 4(b)): GPT-4o leads with 90% (unidiomatic) and 84% (idiomatic) SR, while DeepSeek-R1 follows with 86% and 84%, showing strong performance across both steps. Claude 3.5 also comes with 86% and 84%, performing better than in the other dataset. Llama 3.3 exhibits a significant drops (78% and 73%) in both steps, while Gemini 2.0 maintains stable performance (80% and 77%).
The results show that all models, except DeepSeek-R1, experience a drop in success rate on TransCoder-IR compared to CodeNet. This is due to that TransCoder-IR contains more complex C code with advanced logic and array manipulation compared to CodeNet. However, DeepSeek-R1 maintains consistent performance across both datasets, suggesting that models with reasoning capabilities may be better for handling complex code logic and data manipulation.
6.1.2 Failure Analysis
Table 2: Failure reason categories in TransCoder-IR dataset
Category Description
R1 Memory safety violations in array operations due to improper bounds checking
R2 Mismatched data type translations
R3 Incorrect array sizing and memory layout translations
R4 Incorrect string representation conversion between C and Rust
R5 Failure to handle C’s undefined behavior with Rust’s safety mechanisms
R6 Use of C-specific functions in Rust without proper Rust unsafe wrapper
Cases where SACTOR fails offer valuable insights into areas that require refinement. For each failure case in the two datasets, we conduct an analysis to determine the primary cause of translation failure. This process involves leveraging DeepSeek-R1 to identify potential reasons (prompts available in Appendix F.5), followed by manual verification to ensure correctness. We only focus on the translation process from C to unidiomatic Rust because: (1) it is the most challenging step, and (2) it can better reflect the model’s ability to fit the syntactic and semantic differences between the two languages. Table 2 and 3 summarize the failure reasons, and Figure 5 and 6 illustrate failure reasons (FRs) across models.
(1) TransCoder-IR (Table 2, Figure 5): Based on the analysis, we observe that different models exhibit varying failure reasons. Claude 3.5 shows a particularly high incidence of string representation conversion errors (R4), with 24 out of 45 total failures in the unidiomatic translation step. In contrast, GPT-4o has only 1 out of 17 failures in this category. Llama 3.3 demonstrates consistent challenges with both R4 (string handling) and R6 (using C-specific functions without proper unsafe wrappers), with 10 files for each category. GPT-4o shows a more balanced distribution of errors, with its highest count in R3 (incorrect array sizing and memory layout translations). All models except GPT-4o struggle with string handling (R4) to varying degrees, suggesting this is one of the most challenging aspects of the translation process. DeepSeek-R1 has the fewest overall errors across categories, with failures only in R1 (1 file), R3 (2 files), and R4 (3 files), while completely avoiding errors in R2, R5, and R6.
Figure 5: Failure reasons across different LLM models for the TransCoder-IR dataset.
(2) Project CodeNet (Table 3, Figure 6): Similar to the TransCoder-IR dataset, we also observe that different models in Project CodeNet demonstrate varying failure reasons. C-to-Rust code translation challenges in the CodeNet dataset. Most notably, S6 (mismatched data type translations) presents a significant barrier for Llama 3.3 and Gemini 2.0 (7 files each), while GPT-4o and Claude 3.5 completely avoid this issue. Input argument handling (S1) and format string mistranslations (S3) emerge as common challenges across all models in CodeNet, suggesting fundamental difficulties in translating these language features regardless of model architecture. Only Llama 3.3 and DeepSeek-R1 encounter control flow translation failures (S7), with 2 files each. S4 (random number generation) and S5 (mutable global state variables) are unable to be translated by SACTOR because the current SACTOR implementation does not support these features.
Compared to the results in TransCoder-IR, string representation conversion (R4 in TransCoder-IR, S3 in CodeNet) remains a consistent challenge across both datasets for all models, though the issue is significantly more severe in TransCoder-IR, particularly for Claude 3.5 (24 files). This also suggests that reasoning models like DeepSeek-R1 are better at handling complex code logic and string/array manipulation, as they exhibit fewer failures in these areas, demonstrating the potential of reasoning models to address complex translation tasks.
Table 3: Failure reason categories in Project CodeNet dataset
Category Description
S1 Improper translation of command-line argument handling or try to fix wrong command-line argument handling
S2 Function naming mismatches between C and Rust
S3 Format string directive mistranslation causing output inconsistencies
S4 Original code contains random number generation
S5 SACTOR unable to translate mutable global state variables
S6 Mismatched data type translations
S7 Incorrect control flow or loop boundary condition translations
Figure 6: Failure reasons across different LLM models for the Project CodeNet dataset.
6.2 SACTOR Cost
Table 4: Average Cost Comparison of Different LLMs Across Two Datasets. The color intensity represents the relative cost of each metric for each dataset.
LLM Dataset Tokens Avg. Queries
Claude 3.5 TransCoder-IR 4595.33 5.15
CodeNet 3080.28 3.15
Gemini 2.0 TransCoder-IR 3343.12 4.24
CodeNet 2209.38 2.39
Llama 3.3 TransCoder-IR 6265.13 6.13
CodeNet 3035.43 3.06
GPT-4o TransCoder-IR 2651.21 4.24
CodeNet 2565.36 2.95
DeepSeek-R1 TransCoder-IR 17895.52 4.77
CodeNet 13592.61 3.11
To understand costs, we only analyze programs that successfully generate idiomatic Rust code, excluding failed attempts (as they always reach the maximum retry limit and do not contribute meaningfully to the cost analysis). We evaluate the combined cost of both translation phases to assess overall efficiency. Table 4 compares the average cost of different LLMs across two datasets, measured in token usage and query count per successful idiomatic Rust translation as mentioned in Section 5.2.
Results: Gemini 2.0 and GPT-4o are the most efficient models, requiring the fewest tokens and queries. GPT-4o maintains a low token cost (2651.21 on TransCoder-IR, 2565.36 on CodeNet) with 4.24 and 2.95 average queries, respectively. Gemini 2.0 is similarly efficient, especially on CodeNet, with the lowest token usage (2209.38) and requiring only 2.39 queries on average. Claude 3.5, despite its strong performance on CodeNet, incurs higher costs on TransCoder-IR (4595.33 tokens, 5.15 queries), likely due to additional translation steps. Llama 3.3 is the least efficient in non-thinking model (GPT-4o, Claude 3.5, Gemini 2.0), consuming the most tokens (6265.13 on TransCoder-IR) and requiring the highest number of queries (6.13), indicating significant resource demands.
As a reasoning model, DeepSeek-R1 consumes significantly more tokens (17,895.52 vs. 13,592.61) than non-reasoning models—5-7 times higher than GPT-4o—despite having a similar average query count (4.77 vs. 3.11) for generating idiomatic Rust code. This high token usage comes from the “reasoning process” required before code generation.
6.3 Measuring Idiomaticity
We compare our approach with four baselines: C2Rust (Immunant, 2020), Crown (Zhang et al., 2023), C2SaferRust (Nitin et al., 2025) and Vert (Yang et al., 2024). Of these baselines, C2Rust is the most versatile [^3] Versatility refers to an approach’s applicability to diverse C programs., supporting most C programs, while Crown is also broad but lacks support for some language features. C2SaferRust focuses on refining the unsafe code produced by C2Rust, allowing it to handle a wide range of C programs. In contrast, Vert targets a specific subset of simpler C programs. We assess the idiomaticity of Rust code generated by C2Rust, Crown, and C2SaferRust on both datasets. Since Vert produced Rust code only for TransCoder-IR, we evaluate it solely on this dataset.
Figure 7: Distribution of Total Clippy Issues (Warnings + Errors) across Different Method
Results: Figure 7 presents the lint alert count (sum up of Clippy warnings and errors count for a single program) across all approaches. C2Rust consistently exhibits high Clippy issues, and Crown shows little improvement over C2Rust, indicating both struggle to generate idiomatic Rust. C2SaferRust reduces Clippy issues, but it still retains a significant number of warnings and errors. Notably, even the unidiomatic output of SACTOR surpasses all of these 3. This underscores the advantage of LLMs over rule-based methods. While Vert improves idiomaticity, SACTOR’s idiomatic phase yields fewer Clippy issues, outperforming some existing LLM-based approaches.
Table 5: Unsafe Code Statistics. UF denotes Unsafe Free and AU denotes Avg. Unsafe
Method Dataset SR (%) UF (%) AU (%)
C2Rust TransCoder-IR 100 0 100
CodeNet 100 0 75.9
Crown TransCoder-IR 100 0 100
CodeNet 100 0 75.9
C2SaferRust TransCoder-IR 90 45.6 10.8
CodeNet 93 0 75.8
Vert TransCoder-IR 92 95.7 1.6
SACTOR (Unid.) TransCoder-IR 83 3.6 91.7
CodeNet 90 1.1 42.7
SACTOR (Idiom.) TransCoder-IR 78 100 0
CodeNet 84 100 0
Table 5 summarizes unsafe code statistics. Unsafe-Free indicates the percentage of programs without unsafe code, while Avg. Unsafe represents the average proportion of unsafe code across all translations. C2Rust and Crown generate unsafe code in all programs with a high average unsafe percentage. C2SaferRust has the ability to reduce unsafe code and able to generate unsafe-free programs in some cases (45.6% in TransCoder-IR), but it can not sufficiently reduce the unsafe uses in the CodeNet dataset. Vert has a higher success rate than SACTOR but occasionally introduces unsafe code. SACTOR’s unidiomatic phase retains C semantics, leading to a high unsafe percentage. However, its idiomatic phase eliminates all unsafe code, achieving a 100% Unsafe-Free rate, outperforming all baselines.
6.4 Complex Code-bases
We evaluate the generalization of our approach to complex code-bases by two case studies: the avl_tree and the urlparser; more details are in Appendix D. For both two case studies, we use the GPT-4o model to generate Rust code from the C code. In summary, SACTOR successfully translates the entire C project into unidiomatic Rust, achieving 10/23 function and 4/5 data type conversions. Failures stem from the LLM’s difficulty in generating correct test harnesses for verification.
(a) avl_tree Attempts Distribution
(b) urlparser Attempts Distribution
Figure 8: Success rates (SR) across different LLM models for the TransCoder-IR and CodeNet datasets. SR 1-6 represent the number of attempts made to achieve a successful translation in idiomatic translation.
Case Study 1: avl_tree. The avl_tree project consists of two parts: avl_data, which provides helper functions for data management, and avl_bf, which implements the AVL tree. We successfully generate idiomatic Rust for avl_data, passing verification tests. However, for avl_bf, we only produce unidiomatic Rust, though it still passes verification.
Figure 8(a) shows the generation attempts. Failures stem from Rust compilation errors, which are used as feedback to refine subsequent attempts. Within 4 attempts per function, we achieve idiomatic Rust for avl_data and unidiomatic Rust for avl_bf.
The key challenge in generating idiomatic Rust for avl_bf lies in handling function pointers with void types:
int (*compare)(const void *, const void *);
In unidiomatic Rust, this translates directly using raw pointers:
Option<unsafe extern "C" fn(*const c_void, *const c_void) -> c_int>
For idiomatic Rust, raw pointers should be replaced with generics:
Option<Box<dyn Fn(&T, &T) -> i32>>
where T represents the AVL tree’s data type. However, since our verification tests rely on FFI compatibility, maintaining the C interface requires raw pointers. Rust’s generics are determined at compile time, making it impossible to seamlessly convert them to raw pointers. This constraint prevents our approach from producing idiomatic Rust for avl_bf.
Case Study 2: urlparser. The urlparser project is a simple C-based URL parser. We successfully generate unidiomatic Rust code that passes verification, as shown in Figure 8(b).
For idiomatic Rust, we successfully translate 10 of 23 functions and 4 of 5 data types (enums, structures, and constants). However, the key challenge for the remaining functions is constructing a correct harness function to verify the generated code. As discussed in § 4.3.2, idiomatic Rust functions require a harness to convert C data structures to Rust and transfer outputs back. While the translated Rust code may be correct, the inability to generate a proper harness prevents verification. Without verification, we cannot confirm the correctness of these functions.
7 Conclusion
Translating C to Rust enhances memory safety but is error-prone and often unidiomatic. While LLMs improve translation, they lack correctness guarantees and struggle with semantic differences. SACTOR tackles this with a two-step approach, preserving semantics first, then refining for Rust conventions. Leveraging static analysis and FFI-based validation, it outperforms existing methods, with GPT-4o reaching 80% correctness and DeepSeek-R1 reaching 93% correctness. However, key challenges remain: ensuring correctness, handling complex C features, scaling, interoperability, and efficiency. Some examples of prompts used in SACTOR are shown in Appendix F.
Appendix A A Non-Exhaustive List of Differences Between C and Rust
Here, we present a non-exhaustive list of differences between C and Rust in Table 6, highlighting the key features that make translating code from C to Rust challenging. While the list is not comprehensive, it provides insights into the fundamental distinctions between the two languages, which can help developers understand the challenges of migrating C code to Rust.
Table 6: Key Differences Between C and Rust
Feature C Rust
Memory Management Manual (through malloc/free) Automatic (through ownership and borrowing)
Pointers Raw pointers like *p Safe references like &p/&mut p, Box and Rc
Lifetime Management Manual freeing of memory Lifetime annotations and borrow checker
Error Handling Error codes and manual checks Explicit handling with Result and Option types
Null Safety Null pointers allowed (e.g., NULL) No null pointers; uses Option for nullable values
Concurrency No built-in protections for data races Enforces safe concurrency with ownership rules
Type Conversion Implicit conversions allowed and common Strongly typed; no implicit conversions
Standard Library C stand library with direct system calls Rust standard library with utilities for strings, collections, and I/O
Language Features Procedure-oriented with minimal abstractions Modern features like pattern matching, generics, and traits
Appendix B Algorithm for Task Division
The task division algorithm is used to determine the order in which the items should be translated. The algorithm is shown in Algorithm 1.
Algorithm 1 Translation Task Order Determination
1:$L_{i}$: List of items to be translated
2:$dep(a)$: Function to get dependencies of item $a$
3:$L_{sorted}$: List of groups resolving dependencies
4:$L_{sorted}\leftarrow\emptyset$triangleright Empty list
5:while $|L_{sorted}|<|L_{i}|$ do
6:   $L_{processed}\leftarrow\emptyset$
7:   for $a\in L_{i}$ do
8:      if $a\notin L_{processed}$ and $dep(a)\subseteq L_{processed}$ then
9:         $L_{sorted}\leftarrow L_{sorted}+a$triangleright Add to sorted list
10:         $L_{processed}\leftarrow L_{processed}\cup a$
11:      end if
12:   end for
13:   if $L_{processed}=\emptyset$ then
14:      $L_{circular}\leftarrow DFS(L_{i},dep)$triangleright Circular dependencies
15:      $L_{sorted}\leftarrow L_{sorted}+L_{circular}$triangleright Add a group to sorted list
16:   end if
17:end while
18:return $L_{sorted}$
In the algorithm, $L_{i}$ is the list of items to be translated, and $dep(a)$ is a function that returns the dependencies of item $a$. The algorithm returns a list $L_{sorted}$ that contains the items in the order in which they should be translated. $DFS(L_{i},dep)$ is a depth-first search function that returns a list of items involved in a circular dependency. It begins by collecting all items (e.g., functions, structs) to be translated and their respective dependencies (in both functions and data types). Items with no unresolved dependencies are pushed into the translation order list first, and other items will remove them from their dependencies list. This process continues until all items are pushed into the list, or circular dependencies are detected. If circular dependencies are detected, we resolve them through a depth-first search strategy, ensuring that all items involved in a circular dependency are grouped together and handled as a single unit.
Appendix C Equivalence Testing Details in Prior Literature
C.1 Symbolic Execution-Based Equivalence
Symbolic execution explores all potential execution paths of a program by using symbolic inputs to generate constraints [King, 1976, Baldoni et al., 2018, Coward, 1988]. While theoretically powerful, this method is impractical for verifying C-to-Rust equivalence due to differences in language features. For instance, Rust’s RAII (Resource Acquisition Is Initialization) pattern automatically inserts destructors for memory management, while C relies on explicit malloc and free calls. These differences cause mismatches in compiled code, making it difficult for symbolic execution engines to prove equivalence. Additionally, Rust’s compiler adds safety checks (e.g., array boundary checks), which further complicate equivalence verification.
C.2 Fuzz Testing-Based Equivalence
Fuzz testing generates random or mutated inputs to test whether program outputs match expected results [Zhu et al., 2022, Miller et al., 1990, Liang et al., 2018]. While more practical than symbolic execution, fuzz testing faces challenges in constructing meaningful inputs for real-world programs. For example, testing a URL parsing function requires generating valid URLs with specific formats, which is non-trivial. For large C programs, this difficulty scales, making it infeasible to produce high-quality test cases for every translated Rust function.
Appendix D Dataset Details
D.1 TransCoder-IR Dataset [Szafraniec et al., 2023]
The TransCoder-IR dataset is used to evaluate the TransCoder-IR model and consists of solutions to coding challenges in various programming languages. For evaluation, we focus on the 698 C programs available in this dataset. First, we filter out programs that already have corresponding Rust code. Several C programs in the dataset contain bugs, which are removed by checking their ability to compile. We then use valgrind to identify and discard programs with memory errors during the end-to-end tests. Finally, we select 100 programs with the most lines of code for our experiments.
D.2 Project CodeNet [Puri et al., 2021]
Project CodeNet is a large-scale dataset for code understanding and translation, containing 14 million code samples in over 50 programming languages collected from online judge websites. From this dataset, which includes more than 750,000 C programs, we target only those that accept external input. Specifically, we filter programs using argc and argv, which process input from the command line. As the end-to-end tests are not available for this dataset, we develop the SACTOR test generator to automatically generate end-to-end tests for these programs based on the source code. For evaluation, we select 200 programs and refine the dataset to include 100 programs that successfully generate end-to-end tests.
D.3 Real-World Projects
For § 6.4, we use two real-world projects for evaluation: avl-tree [^4] https://github.com/xieqing/avl-tree and urlparser [^5] https://github.com/jwerle/url.h. Both projects are written in C and have some non-trivial C code features. While earlier works may have used different versions, we selected these projects because they include end-to-end tests, enabling us to evaluate the correctness of the translated code. To make the projects fit the input requirements, we use the cpp preprocessor to expand macros and combine the entire project into a single file. The avl-tree project contains 12 end-to-end tests to test the different functionalities of the AVL tree implementation. The urlparser project contains 3 end-to-end tests to test the different functionalities of the URL, we manually create 7 additional end-to-end tests to test the different functionalities of the URL parser.
Appendix E LLM Configurations
Table 7 shows our configurations for different LLMs in evaluation. All other hyper-parameters, like Top-P or Top-K, are set as the model’s default values.
Table 7: Configurations of Different LLMs in Evaluation
Model Version Temperature Hosting Platform
GPT-4o gpt-4o-latest 1 AzureOpenAI API
Claude 3.5 Sonnet claude-3-5-sonnet-20241022 1 Anthropic API
Gemini 2.0 Flash gemini-2.0-flash-exp default Google Cloud API
Llama 3.3 Instruct 70B Llama 3.3 Instruct 70B Q4 0.8 Single H100 GPU
DeepSeek-R1 DeepSeek-R1 671B 1 DeepSeek API
Appendix F Examples of Prompts Used in SACTOR
The following prompts are used to guide the LLM in C-to-Rust translation and verification tasks. The prompts may slightly vary to accommodate different translation task, as SACTOR leverages static analysis to fetch the necessary information for the LLM.
F.1 Unidiomatic Translation
Figure 9 shows the prompt for translating unidiomatic C code to Rust.
Translate the following C function to Rust. Try to keep the **equivalence** as much as possible.
libc will be included as the **only** dependency you can use. To keep the equivalence, you can use unsafe if you want.
The function is:
‘‘‘c
{C_FUNCTION}
// Specific for main function
The function is the main function, which is the entry point of the program. The function signature should be: pub fn main() -> ()‘.
For return 0;‘, you can directly return;‘ in Rust or ignore it if its the last statement.
For other return values, you can use std::process::exit()‘ to return the value.
For argc and argv‘, you can use std::env::args()‘ to get the arguments.
The function uses some of the following stdio file descriptors: stdin. Which will be included as
‘‘‘rust
extern "C" {
static mut stdin: *mut libc::FILE;
You should **NOT** include them in your translation, as the system will automatically include them.
The function uses the following functions, which are already translated as (you should **NOT** include them in your translation, as the system will automatically include them):
‘‘‘rust
{DEPENDENCIES}
Output the translated function into this format (wrap with the following tags):
----FUNCTION----
‘‘‘rust
// Your translated function here
----END FUNCTION----
Figure 9: Unidiomatic Translation Prompt
F.2 Unidiomatic Translation with Feedback
Figure 10 shows the prompt for translating unidiomatic C code to Rust with feedback from the previous incorrect translation and error message.
Translate the following C function to Rust. Try to keep the **equivalence** as much as possible.
libc will be included as the **only** dependency you can use. To keep the equivalence, you can use unsafe if you want.
The function is:
‘‘‘c
{C_FUNCTION}
// Specific for main function
The function is the main function, which is the entry point of the program. The function signature should be: pub fn main() -> ()‘.
For return 0;‘, you can directly return;‘ in Rust or ignore it if its the last statement.
For other return values, you can use std::process::exit()‘ to return the value.
For argc and argv‘, you can use std::env::args()‘ to get the arguments.
The function uses some of the following stdio file descriptors: stdin. Which will be included as
‘‘‘rust
extern "C" {
static mut stdin: *mut libc::FILE;
You should **NOT** include them in your translation, as the system will automatically include them.
The function uses the following functions, which are already translated as (you should **NOT** include them in your translation, as the system will automatically include them):
‘‘‘rust
fn atoi (str : * const c_char) -> c_int;
Output the translated function into this format (wrap with the following tags):
----FUNCTION----
‘‘‘rust
// Your translated function here
----END FUNCTION----
Lastly, the function is translated as:
‘‘‘rust
{COUNTER_EXAMPLE}
It failed to compile with the following error message:
{ERROR_MESSAGE}
Analyzing the error messages, think about the possible reasons, and try to avoid this error.
Figure 10: Unidiomatic Translation with Feedback Prompt
F.3 Idiomatic Translation
Figure 11 shows the prompt for translating unidiomatic Rust code to idiomatic Rust. Crown is used to hint the LLM about the ownership, mutability, and fatness of pointers.
Translate the following unidiomatic Rust function into idiomatic Rust. Try to remove all the unsafe blocks and only use the safe Rust code or use the unsafe blocks only when necessary.
Before translating, analyze the unsafe blocks one by one and how to convert them into safe Rust code.
**libc may not be provided in the idiomatic code, so try to avoid using libc functions and types, and avoid using std::ffi module.**
‘‘‘rust
{RUST_FUNCTION}
"Crown" is a pointer analysis tool that can help to identify the ownership, mutability and fatness of pointers. Following are the possible annotations for pointers:
fatness:
- Ptr‘: Single pointer
- Arr‘: Pointer is an array
mutability:
- Mut‘: Mutable pointer
- Imm‘: Immutable pointer
ownership:
- Owning‘: Owns the pointer
- Transient‘: Not owns the pointer
The following is the output of Crown for this function:
{CROWN_RESULT}
Analyze the Crown output firstly, then translate the pointers in function arguments and return values with the help of the Crown output.
Try to avoid using pointers in the function arguments and return values if possible.
Output the translated function into this format (wrap with the following tags):
----FUNCTION----
‘‘‘rust
// Your translated function here
----END FUNCTION----
Figure 11: Idiomatic Translation Prompt
F.4 Idiomatic Verification
Idiomatic verification is the process of verifying the correctness of the translated idiomatic Rust code by generating a test harness. The prompt for idiomatic verification is shown in Figure 12.
This is the idiomatic translation of Rust code from C, the function signature is
‘‘‘rust
{FUNCTION_NAME}_idiomatic
This is the unidiomatic translation of Rust code from C, the function signature is
‘‘‘rust
{FUNCTION_DEPENDENCIES}
Generate the harness for the function atoi_idiomatic with the following code pattern so that it can be tested:
Finish all the TODOs.
You should **NOT** add any dummy implementation of the function or structs, as it will be provided by the verifier:
‘‘‘rust
// TODO: add necessary uses here
// Dont add the definitions of any other functions and structs, they will be provided by the system
fn {FUNCTION_NAME} ({UNIDIOMATIC_ARGS}) -> {UNIDIOMATIC_RETURN} {
// TODO: Add code here to Convert the input to the idiomatic format
let result = fn {FUNCTION_NAME}_idiomatic({IDIOMATIC_ARGS}) -> {IDIOMATIC_RETURN}; // Call the idiomatic function
// TODO: Add code here to Convert the result back to the original format
remove all the TODOs and replace them with the necessary code.
Output the translated function into this format (wrap with the following tags):
----FUNCTION----
‘‘‘rust
// Your translated function here
----END FUNCTION----
Figure 12: Idiomatic Verification Prompt
F.5 Failure Reason Analysis
Figure 13 shows the prompt for analyzing the reasons for the failure of the translation.
Given the following C code:
‘‘‘c
{original_code}
The following code is generated by a tool that translates C code to Rust code. The tool has a bug that causes it to generate incorrect Rust code. The bug is related to the following error message:
‘‘‘json
{json_data}
Please analyze the error message and provide a reason why the tool generated incorrect Rust code.
1. Append a new reason to the list of reasons.
2. Select a reason from the list of reasons that best describes the error message.
Please provide a reason why the tool generated incorrect Rust code **FUNDAMENTALLY**.
List of reasons:
{all_current_reasons}
Please provide the analysis output in the following format:
‘‘‘json
"action": "append", // or "select" to select a reason from the list of reasons
"reason": "Format string differences between C and Rust", // the reason for the error message, if action is "append"
"selection": 1 // the index of the reason from the list of reasons, if action is "select"
// "reason" and "selection" are mutually exclusive, you should only provide one of them
Please **make sure** to provide a general reason that can be applied to multiple cases, not a specific reason that only applies to the current case.
Please provide a reason why the tool generated incorrect Rust code **FUNDAMENTALLY** (NOTE that the reason of first failure is always NOT the fundamental reason).
Figure 13: Failure Reason Analysis Prompt